OpenStudy (jokersmart):

Check my answer? Help please! Choose all that apply. Dennis is entering a contest to guess the number of marbles inside a cylinder. He wants to find the volume of the marbles and the volume of the cylinder to estimate how many marbles might be in the cylinder. Which formulas does he need to make a good guess? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.)

OpenStudy (jokersmart):

\[A. V=lwh\] \[B. V=\frac{ 4 }{ 3 }\pi r ^{3}\] \[C. V=\pi r ^{2}h\] \[D. V=\frac{ 1 }{ 3 }\pi r ^{2} h\]

OpenStudy (jokersmart):

I say B and C

OpenStudy (jokersmart):

@Catlover5925 Its kinda hard but I appreciate your help

OpenStudy (jokersmart):

I understand if you cant though

OpenStudy (catlover5925):

yes i agree!

OpenStudy (jokersmart):

Awesome, I'll see what I get! Thanks! And i hope everything goes well with you :)

OpenStudy (catlover5925):

thanks...let me know if its right...i dont like it when i do something wrong

OpenStudy (jokersmart):

You did it right as i can remember